An advertising agent goes on a vacation to three countries, India, New Zealand and Australia. He spends three times the number of days in New Zealand than he does in Australia and five days longer in India than in Australia. If 40 was a vacation days how many did he spend In Australia?


1
Expert's answer
2019-09-12T09:29:52-0400

Let i be the number of day in India, n - in New Zealand, a - in Australia.

n=3a

i=a+5

i+n+a=40

So, a+5+3a+a=40

5a=35

a=7